OpenStudy (anonymous):

rational numbers unclude whole numbers , integers , fractions and some decimals ? is it true or false thank you

OpenStudy (anonymous):

The rational number is a real number that can be written as a fraction. So yes decimals like 0.5 would work because you could write it as: 1/2 \[\pi \] could not be a rational number because it goes on forever and it doesn't repeat! 3.141.... etc However, numbers do have a recurring decimal point like 1/3 = 0.33333 it is rational because you can still write it as a fraction. ;p

OpenStudy (unklerhaukus):

whole number can be written as fraction where the denominator is 1 eg 4 = 4/1